OpenStudy (anonymous):

derivative of 36x^2 + 100y^2 - 108x + 500y + 589

OpenStudy (anonymous):

use the power rule

OpenStudy (anonymous):

36(2) x + 200y -108 + 500 72x + 200y + 392

OpenStudy (anonymous):

got - 72x + 108 / 200y + 500

OpenStudy (anonymous):

am i right?

zepdrix (zepdrix):

Is your derivative being taken with respect to x or y?

OpenStudy (anonymous):

dy/dx

zepdrix (zepdrix):

So any time we take the derivative of a non-x term we need to apply the chain rule and multiply by dy/dx. So we start by applying the power rule to every term, and then the chain rule to all the y terms.\[\Large \frac{d}{dx}(36x^2 + 100y^2 - 108x + 500y + 589 )\] \[\Large = 36(2x)+100(2y)\frac{dy}{dx}-108(1)+500(1)\frac{dy}{dx}+0\]

OpenStudy (anonymous):

dy/dx is - 72x + 108 / 200y + 500?

zepdrix (zepdrix):

Yah that looks right, good job! :) dy/dx = (-72x + 108) / (200y + 500)

OpenStudy (anonymous):

thanks thanks